Project TBASERDN031029: Goal: investigate X-band receiver Setup: LCP --> ODM6 --> CM9 --> SF1 --> J1 Bank A, 1 CM10 --> SF2 --> J3 Bank B, 1 CM11 --> SF3 --> J5 Bank C, 1 RCP --> ODM8 --> CM13 --> SF5 --> J2 Bank A, 2 CM14 --> SF6 --> J4 Bank B, 2 CM15 --> SF7 --> J6 Bank C, 2 Due to missing G8, have to get to SF4 and SF8 via OR1, 2, CM4, 8. LCP --> ODM1 --> CM4 --> SF4 --> J7 Bank D, 1 RCP --> ODM3 --> CM8 --> SF8 --> J8 Bank D, 2 Tune: Bank A,B,C, D to 9000MHz Bank LO2 ---- ------ A 12300 B 12300 C 12300 D 12300 ----------- setup script: /users/rnorrod/gbtgui config file: /home/groups/baseline/Projects/TBASERDN031023/Rx8_10acs.cfg Scan Source Proc Mode Comments ----- -------- -------- ---------- ---------------------------------------------------- 2-3 NGC 7027 xonoff samplers balanced, on IF noise source!! 4-5 NGC 7027 xonoff samplers balanced Conv Atten: CM9 8.25 SF1 CM10 7.25 SF2 CM11 11.125 SF3 CM4 10.5 SF4 CM13 6.375 SF5 CM14 7.25 SF6 CM15 8.875 SF7 CM8 3.375 SF8 IF Rack Atten: ODM6 21 ODM8 22 6 NGC 7027 track samplers not rebalanced samplers 1 and 2 clipping levels reset for -20 dBm optimum input level instead of -17 dBm normally Conv Atten: CM9 8.25 SF1 <- CM10 7.25 SF2 CM11 11.125 SF3 CM4 10.5 SF4 CM13 6.375 SF5 <- CM14 7.25 SF6 CM15 8.875 SF7 CM8 3.375 SF8 IF Rack Atten: ODM6 21 ODM8 22 7-8 NGC 7027 xonoff samplers not rebalanced but changed CM attenuators by +3 dB in CM9 and CM13 samplers 1 and 2 clipping levels reset for -20 dBm optimum input level instead of -17 dBm normally Conv Atten: CM9 11.25 SF1 <- CM10 7.25 SF2 CM11 11.125 SF3 CM4 10.5 SF4 CM13 9.375 SF5 <- CM14 7.25 SF6 CM15 8.875 SF7 CM8 3.375 SF8 IF Rack Atten: ODM6 21 ODM8 22 9-10 NGC 7027 xonoff samplers not rebalanced but changed CM attenuators by -3 dB in CM9 and CM13 samplers 1 and 2 clipping levels reset for -14 dBm optimum input level instead of -17 dBm normally, RA pointing off by about 4 sec of time Conv Atten: CM9 5.25 SF1 <- CM10 7.25 SF2 CM11 11.125 SF3 CM4 10.5 SF4 CM13 3.375 SF5 <- CM14 7.25 SF6 CM15 8.875 SF7 CM8 3.375 SF8 IF Rack Atten: ODM6 21 ODM8 22 11-12 NGC 7027 xonoff same as 9-10 but with pointing trimmed up 13-14 NGC 7027 xonoff sampler 1 and 2 back to niminal level settings. SF filters set to 200 MHz BW while spectrometer running at 800 MHz BW. All levels rebalanced but inputs inputs to 5 and 6 mis-set,